OpenStudy (anonymous):

Evaluate indefinite integral sec(3x)tan(3x) dx

OpenStudy (zarkon):

\[\frac{d}{dx}\sec(3x)=\sec(3x)\tan(3x)3\]

OpenStudy (jhannybean):

let u=3x, du = 3dx -> du/3 = dx \[\large \frac{1}{3}\int\limits \sec(u)\tan(u)du =?\]

OpenStudy (anonymous):

thanx!!

OpenStudy (anonymous):